OpenStudy (anonymous):

Find a quadratic function that has a y-intercept of y=4 and is increasing on the point (3,28).

OpenStudy (alekos):

have you tried to solve this?

OpenStudy (anonymous):

no i cant do it

OpenStudy (anonymous):

i think its f(x)=x^2-4

OpenStudy (alekos):

no, if you let x=0 which is the y intercept then you get y=-4 let me assist

OpenStudy (alekos):

the quadratic has to be on the form y=ax^2+bx+c when x=0 the y intercept is c so what do you think c should be?

OpenStudy (anonymous):

i dont know

OpenStudy (alekos):

just substitute x=0 into the general quadratic equn and you will get a y value

OpenStudy (alekos):

and you already know that the y-intercept has to be 4 i.e. when x=0, y=4

OpenStudy (alekos):

so what should c equal?